From the circuit shown below,

a M 10 V 30 il 222 i2 722 892 i3 12 V W

(a) use KCL and KVL to solve for the three currents.

(b) With the given data below and the two currents flowing clockwise, determine both currents using mesh analysis and the voltages across each resistor.

r1= 3 Ohms

r2= 7 Ohms

r3= 2 Ohms

r4= 8 Ohms
Answer #1

ΤΟΥ a LI s (a) By KCL,. Vat 10 M 31 an + Vat Va-12 20 8 2 7 + 3 81 mu 2.3 12 V ov, Yal? + + + + + 2) = (1 - 1) . 0.725va = -010V. + 30 312 B) By 0 (2) FO Vat By KVL in loop di 10 + 2 (5-५) + 81320 5 +i -i ++i320 14 =_51, +5 ------ 22 7/- 8.2 + 121 ४e By KVL in loop , 12 1 12 + 1014 + 2(14-1₂)=0 6+514 + 14 13 il Or, 6+614 -13 we know; in a 513 + 5 6+ 6( 5i3+5) - 3-0 oy, 6+